1, 1, 2, 4, 12,___, 108 find the value in the blank
Found 2 solutions by FrankM, Edwin McCravy:
Answer by FrankM(1040)

36
1*1=1
1*2=2
2*2=4
4*3=12
12*3=36
36*3=108
The next number would be
108*4=432
and we'd multiply by 4 4times, then mover to 5.

Answer by Edwin McCravy(20056)

1, 1, 2, 4, 12,___, 108
Start with 1.

Use 1 as a multiplier of the previous term for the next 1 term.

1x1 = 1

Use 2 as a multiplier of the previous term for the next 2 terms.

1x2 = 2
2x2 = 4

Use 3 as a multiplier of the previous term for the next 3 terms.

4x3 = 12
12x3 = 36       <-- (that's the answer, but let's keep going a little more.)
36x3 = 108

Use 4 as a multiplier of the previous term for the next 4 terms.

108x4 = 432
432x4 = 1728
1728x4 = 6912
6912x4 = 27648

Edwin